OpenStudy (anonymous):

The surface area of a cone is 16 pi inches squared. The radius is 3 inches. What is the slant height?

OpenStudy (jdoe0001):

hmmm

OpenStudy (jdoe0001):

am I wrong by using instead \(\bf \textit{area of a cone } = \pi r \sqrt{r^2+h^2} + \pi r^2\\ \color{blue}{h = \textit{height of the cone}\\ \sqrt{r^2+h^2} = \textit{slant height}}\\ \textit{area of a cone } = \pi r (\textit{slant height})+ \pi r^2 \implies \pi r ((\textit{slant height})+r)\\ 16 \pi = \pi 3 ((\textit{slant height})+3)\)

OpenStudy (jdoe0001):